OpenStudy (anonymous):

y^-2/y^2 can someone help me understand how to do this?

OpenStudy (anonymous):

\[\frac{ y^-2}{ y^2 }\]

OpenStudy (anonymous):

\(\Large y^{-2}=\frac{1}{y^2}\) So, \(\Large \frac{y^{-2}}{y^2}=\huge \frac{\frac{1}{y^2}}{y^2}\Large= \frac{1}{y^2\times y^2}=\frac{1}{y^4}\)

OpenStudy (anonymous):

\[y^{-2}=\frac{ 1 }{ y }\] so(1/y^2)X(1/y^2) =\[\frac{ 1 }{ y ^{4} }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

u r welcome